# Efficient Dataset Distillation via Diffusion-Driven Patch Selection for Improved Generalization

### v3

@@ 4.2 Performance Improvements
Distinct from all previous methods, our proposed method achieves a fully unrestricted distillation process and demonstrates SOTA performance. The limitations of existing dataset distillation methods are outlined as follows, and Tab. 3 represents a comprehensive comparison:
  • Large-scale datasets. Optimization-based dataset distillation methods using information matching often suffer from excessive computational overhead, rendering them impractical for large-scale datasets and leading to sub-optimal performance.
@@
  • Zero-shot. Relying on pre-trained classifiers require selecting an appropriate model architecture and training it for the target dataset, the nasty teachers even disrupt the distillation process. Although Minimax attempts to address this issue through the use of diffusion model, it is still constrained by DiT trained on ImageNet-1K.
@@
  • Class-combination. Repeated distillation processes are required when dealing with different class-combination. For example, in ImageNet-1K, a 10-class classification task can result in $C_{1000}^{10}$ possible combinations, leading to significant computational waste.
@@
  • IPC. Similarly, varying IPC settings necessitate multiple distillation, also causing substantial time overhead. While D4M addresses the class-combination issue, it continues to necessitate recalculating different numbers of cluster centers with different IPC settings.
Method Large-scale Zero-shot Class-combination IPC
@@
Teacher Network Student Network
ResNet-18 59.4 62.7 65.4 53.4 60.1 62.1 69.2
ResNet-50 52.2 64.8 66.7 48.9 56.2 51.3 60.1
ResNet-101 50.1 63.7 66.5 46.1 55.5 42.8 50.5
MobileNet-V2 55.2 62.5 65.0 55.3 60.5 57.2 62.3
EfficientNet-B0 45.8 59.1 61.1 45.0 58.4 51.7 60.3
Swin-T 34.4 51.2 52.7 35.1 44.1 51.1 50.2
ViT-B 31.2 46.9 49.3 32.7 40.0 41.3 46.2
4.3 Cross-architecture Generalization
Following previous work, we evaluate the performance of the synthetic dataset using different teacher-student combinations. It can be observed that, when using similar network architectures as teacher-student pairs (e.g., both CNNs), shallower networks generally yield better performance as teacher models, while the larger networks show stronger learning ability and obtain the best performance when the size of synthetic datasets increases. We further compare the performance of the synthetic dataset on extremely different network architectures, e.g., CNNs and ViTs. As a student network, the ViT-based networks leverages its global attention mechanism to attain the comparable performance with more realistic images. However, as a teacher network, ViT-based networks does not have strong ability to teach the student networks, yielding weak performance.
@@ 4.4 Ablation Study
w/o Cluster 42.4 58.1 61.0
w/ Cluster-CLIP 42.1 59.0 60.7
w/ Cluster-DIFT 43.9(+1.5) 59.4(+1.3) 61.8(+0.8)
Clustering strategy.
To investigate the effectiveness of the two main stages of our proposed method, we conduct an ablation study to investigate whether clustering operations enhance the representative of synthetic datasets. Moreover, we conduct experiments to explore the differences between using CLIP features and DIFT during the clustering process. The experimental results as shown in Tab. 5, demonstrates that the performance of the synthetic dataset improves further after incorporating the clustering operation. This indicates that reducing sample randomness enables the model to learn more representative information. Additionally, we found DIFT more effective for achieving clustering operations compared to CLIP feature.
Real Images number.
Although performing the distillation operation once on the entire dataset offers the potential for one-step processing, the substantial time overhead makes this approach difficult to apply in practice. To address this issue, we explore the effect of the number of original images used for key patch extraction on performance. As shown in Fig. 6, we randomly select images as a subset of the original dataset and evaluate the corresponding performance. Given that random selection is an unbiased data sampling method, our proposed achieves performance comparable to others, even when the number of original samples is limited. When 300 images are randomly selected from original dataset as input, the performance is nearly on par with the full dataset, resulting in a 4x speedup for the ImageNet-1K dataset.
Figure 6: Ablation on number of real images used.
@@ 5 Conclusion
Supplementary Material
1 Training Efficiency
For a more comprehensive comparison, we evaluate the synthesis time per image and GPU memory consumption of methods that can scale on large-scale datasets. Since previous works on different settings (e.g., different class number of target dataset) require repeated distillation, we measure the time required to generate a single image on the ImageNet-1K under IPC=100 to ensure the fairness. The experimental results are shown in Tab. 6, our method not only achieves the best results but also maintains training efficiency.
SRe2L Minimax D4M RDED Ours
Time (h) 106 256 52 1 95
Accuracy (%) 52.8 58.8 59.3 60.5 61.8
Table 6: Synthesis time per image, GPU memory consumption and corresponding accuracy on ImageNet-1K under IPC=50.
2 Diffusion time intervals
To calculate the difference in predicted loss corresponding to UNet when using different prompts (i.e., w. and w.o. label text), multiple diffusion time steps $t$ are sampled as time conditions and averaged. Therefore, we investigated whether varying the range of time step values would impact the performance of the synthetic dataset.
As shown in Tab. 7, we selected different minimum and maximum edge as the range for the time steps and evaluated the corresponding performance of the synthetic dataset. It can be observed that the dataset achieves the best performance when using intermediate time steps and the small time steps are essential, which is consistent with the conclusions drawn in recent works [18] [32]. In practical applications, considering the generality of ImageNet-1K, we applied the same range of diffusion time steps to all of its subsets (e.g., ImageWoof).
End
Start 0.1 0.3 0.5 0.7 0.9
0.1 - 58.5 58.8 59.4 59.2
0.3 - - 58.1 58.4 59.0
0.5 - - 57.5 58.1 58.6
0.7 - - - - 57.2
3 Time steps on the computation of DIFT
An important component of our proposed framework involves calculating the DIFT [32] for all patches within the same class and using it as the feature for clustering. The process of calculating DIFT first requires performing a diffusion operation on the input patch at specific time steps. We conduct ablation experiments to investigate the impact of using different time steps to calculate DIFT on the performance of the synthetic dataset.
@@ 4 Generalization Ability
As shown in Tab. 8, our proposed method still demonstrates strong performance on CNNs. Additionally, it can be observed that ResNet18 is the most suitable architecture to be used as a teacher model when the sample size is smaller. However, for ViTs, due to their low-pass filter characteristics, effective learning becomes challenging when the sample size is limited and class-specific features are incomplete. The best performance achieved when using DiT as a student model is only 50% of that of the CNN student model under the same settings.
5 Number of cluster centers
@@
Image Size: 128x128
112x112 42.6 70.2 77.8
Image Size: 256x256
112x112 40.4 67.4 76.8
224x224 44.3 73.5 80.2
Image Size: 512x512
112x112 38.6 65.3 71.4
224x224 43.1 69.5 78.4
448x448 45.1 73.2 81.0
6 Impact of Resizing
When using diffusion to process images, the image size often has a certain degree of impact. On the other hand, when calculating $R(p|c)$ for each patch, we used a fixed-size average pooling kernel to obtain the mean value of the patch. Therefore, we designed an ablation experiment to explore whether different input image sizes and patch sizes significantly affect the performance of the synthetic dataset. In practical applications, we proportionally scale the original image such that its size of shortest edge matches the selected size. For different patch sizes, under the setting of IPC=10, we scale all patches to 112x112 in order to concatenate 4 patches into a 224x224 image. Under the setting of IPC=50, we scale all patches to 224x224.
The experimental results, as shown in Tab. 9, indicate that when the original image size closely matches the standard output size of SD v1.5, better performance are achieved. Additionally, when the size of the cropped patches is closer to the scaled original image size, the image tends to exhibit more complete class-relevant features, leading to better performance. Consider the balance between performance and efficiency, We choose 256x256 as the size of the real images, and 224x224 as the size of the patches. Comparison of visualization with different crop sizes when using 256x256 as the size of the real images are shown in Fig. 9, An excessive size difference between patches and the original image can cause the generated dataset to overly focus on low-frequency features (e.g., noses of dogs), leading to performance degradation.
@@ 6 Impact of Resizing
7 Qualitative Analysis
To better understand the superiority of the datasets generated by our proposed method, we used t-SNE to visualize and analyze datasets produced by different methods. As shown in Fig. 10, it is evident that the dataset generated by RDED [31] is overly concentrated around class centers because it originates from a pre-trained classifier. Additionally, since the pre-trained classifier cannot guarantee accurate classification results, the synthetic dataset from RDED suffers from class mixing issues. Similarly, the Minimax [7], which uses the DiT pre-trained on ImageNet-1K to generate images, also results in datasets that are overly concentrated in class centers. This concentration can be detrimental to model learning when IPC increases.
On the other hand, D4M [30] generates images using SD v1.5, which is entirely unrelated to ImageNet-1K, resulting in datasets with excessive noise and outlier samples. Although this approach can broadly cover the original input distribution, the overly vague classification boundaries hinder its effectiveness in providing useful guidance for dataset distillation tasks. In contrast to all the aforementioned methods, our approach achieves a good balance between diversity and representativeness by using SD v1.5 for localization rather than generation. Furthermore, our method can still leverage diffusion models to generate images, allowing for a mixed dataset of original and generated images, which may lead to improved performance.
8 Visualizations
To facilitate a more intuitive understanding of our methodology, we present the complete synthetic datasets for ImageWoof and ImageNette at IPC=10 as illustrated in Figs. 11 and 12, as well as the partial synthetic dataset at IPC=50, depicted in Figs. 13 and 14. Furthermore, we have visualized all images belonging to the “Golden Retriever” class at IPC=100, as shown in Fig. 15.

### v4

@@ 4.4 Training Efficiency
Distinct from all previous methods, our proposed method achieves a fully unrestricted distillation process and demonstrates SOTA performance. The limitations of existing dataset distillation methods are outlined as follows, and Table 4 represents a comprehensive comparison:
  • Large-scale. information-matching dataset distillation methods are impractical for large-scale datasets due to massive time cost.
@@
  • Zero-shot. Relying on pre-trained classifiers require selecting an appropriate model architecture and training it for the target dataset. Minimax is also constrained by DiT pre-trained on ImageNet-1K.
@@
  • Class. Repeated distillation processes are required when dealing with different class-combination. For example, in ImageNet-1K, a 10-class classification task can result in $C_{1000}^{10}$ possible combinations.
@@
  • IPC. Varying IPC settings necessitate multiple distillation. While D4M addresses the class-combination issue, it necessitates recalculating different prototypes with different IPC settings.
For a more comprehensive comparison, we evaluate the total synthesis time of different methods under challenging compression settings. As shown in Figure 5, our one-step distillation method achieves significant efficiency across various tasks, demonstrating a significant improvement in computational efficiency compared to previous approaches.
Method Large-scale Zero-shot Class IPC
@@ 4.5 Ablation Study
w/o Cluster 43.6 58.1 62.0
w/ Cluster-CLIP 42.1 59.0 62.7
w/ Cluster-DIFT 46.1(+2.5) 61.0(+1.9) 63.4(+1.4)
Clustering Strategy. To investigate the effectiveness of the two main stages of our proposed method, we conduct an ablation study to investigate whether clustering operations enhance the representative of synthetic datasets. Moreover, we conduct experiments to explore the differences between using CLIP features and DIFT during the clustering process. The experimental results as shown in Table 5, demonstrates that the performance of the synthetic dataset improves further after incorporating the clustering operation. This indicates that reducing sample randomness enables the model to learn more representative information. Additionally, we found DIFT more effective for achieving clustering compared to CLIP feature.
Diffusion Time Intervals. To calculate the difference in predicted loss corresponding to UNet when using different prompts (i.e., w. and w.o. label text), multiple diffusion time steps $t$ are sampled as time conditions and averaged. Therefore, we investigated whether varying the range of time step values would impact the performance of the synthetic dataset.
As shown in Tab. 6, we selected different minimum and maximum edge as the range for the time steps and evaluated the corresponding performance of the synthetic dataset. It can be observed that the dataset achieves the best performance when using intermediate time steps and the small time steps are essential, which is consistent with the conclusions drawn in recent works [li2023your] [tang2023emergent]. In practical applications, considering the generality of ImageNet-1K, we applied the same range of diffusion time steps to all of its subsets (e.g., ImageWoof).
End
Start 0.1 0.3 0.5 0.7 0.9
0.1 - 58.9 59.3 61.0 59.7
0.3 - - 58.8 58.9 59.6
0.5 - - 58.1 58.6 59.0
0.7 - - - - 57.9
Time Steps on Computation of DIFT. An important component of our proposed framework involves calculating the DIFT [tang2023emergent] for all patches within the same class and using it as the feature for clustering. The process of calculating DIFT first requires performing a diffusion operation on the input patch at specific time steps. We conduct ablation experiments to investigate the impact of using different time steps to calculate DIFT on the performance of the synthetic dataset.
Figure 7: Ablation study on the number of real images used.
Number of Real Images. Although performing the distillation operation once on the entire dataset offers the potential for one-step processing, the substantial time overhead makes this approach difficult to apply in practice. To address this issue, we explore the effect of the number of original images used for key patch extraction on performance. As shown in Figure 7, we randomly select images as a subset of the original dataset and evaluate the corresponding performance. Given that random selection is an unbiased data sampling method, our proposed achieves performance comparable to others, even when the number of original samples is limited. When 300 images are randomly selected from original dataset as input, the performance is nearly on par with the full dataset, resulting in a $4\times$ speedup.
@@ 5 Conclusion
A More Experimental Results
A.1 Experimental Settings
Datasets and Architectures. To evaluate our proposed method on low-resolution datasets, we conduct experiments on CIFAR-10/100 [krizhevsky2009learning] and TinyImageNet [le2015tiny], with ConvNet [gidaris2018dynamic] serving as the backbone. For high-resolution data, we perform evaluations on ImageNet-1K [deng2009imagenet] and its widely used subsets, including ImageNet-100 [kim2022dataset], ImageNette and ImageWoof [cazenavette2022dataset], utilizing ResNet-18 [he2016deep] as the backbone.
Evaluation Protocol. In line with previous work, we set IPC=10 and 50 for low-resolution datasets, and IPC=10, 50, and 100 for high-resolution datasets to enable more efficient learning. During the training process, we employ a fixed pre-trained classifiers to provide soft labels. Subsequently, five randomly initialized networks are trained on the distilled dataset, with the average performance of these networks evaluated.
Implementation Details. When calculating $M(\mathbf{z}|c)$ of images, we select a time-step range of [0.1, 0.7], whereas for the computation of DIFT, a time-step of 0.16 is employed. Across all IPC settings, we prioritize cropping the patches with resolution of $224\times 224$ and selecting them from the top 10 ranked clusters after clustering. All the experiments were carried out using PyTorch on NVIDIA RTX-3090 GPUs.
A.2 Hard Label Performance
To further investigate the quality of synthetic datasets generated by different distillation methods, we conduct an experiment with one-hot labels.
As shown in Tab. 7, Sre2L exhibit extreme poor performance, Due to the lack of diversity of generated images caused by using only cross-entropy loss. For Minimax and RDED, the synthetic datasets also exhibit concentration of class centers which means the loss of diversity, leading a performance degradation especially on large IPC settings. Despite D4M successfully enhance diversity of synthetic datasets by using a SD trained on LAION, the misalignment phenomenon observed between images and corresponding class labels, leading to a significant performance drop when using hard label. In contrast, our method achieve a balance between representative and diversity which has also been verified using t-SNE visualization, thus a strong performance achieved using only hard label.
Table 7: Hard label performance on ImageNet-1K with different methods.
IPC Random SRe2L Minimax D4M RDED Ours
[…]
10 7.3 $\pm$0.2 1.4 $\pm$ 0.2 8.2 $\pm$ 0.4 4.6 $\pm$ 0.5 11.4 $\pm$ 0.4 10.5 $\pm$ 0.3
50 30.9 $\pm$0.1 3.2 $\pm$ 0.3 31.3 $\pm$ 0.4 17.6 $\pm$ 0.2 30.1 $\pm$ 0.3 33.2 $\pm$ 0.1
100 41.0 $\pm$0.3 4.1 $\pm$ 0.3 40.5 $\pm$ 0.2 27.3 $\pm$ 0.5 38.4 $\pm$ 0.2 42.1 $\pm$ 0.2
@@ A.3 Generalization Ability
As shown in Tab. 8, our proposed method still demonstrates strong performance on CNNs. Additionally, it can be observed that ResNet-18 is the most suitable architecture to be used as a teacher model when the sample size is smaller. However, for ViTs, due to their low-pass filter characteristics, effective learning becomes challenging when the sample size is limited and class-specific features are incomplete. The best performance achieved when using DiT as a student model is only 50% of that of the CNN student model under the same settings.
B Additional Ablation Studies
B.1 Number of Cluster Centers
@@
Image Size: $128\times 128$
$112\times 112$ 42.6 70.2 77.8
Image Size: $256\times 256$
$112\times 112$ 40.4 67.4 76.8
$224\times 224$ 44.5 73.7 80.1
Image Size: $512\times 512$
$112\times 112$ 38.6 65.3 71.4
$224\times 224$ 43.1 69.5 78.4
$448\times 448$ 45.1 73.2 81.0
@@
B.2 Impact of Resizing
When using diffusion to process images, the image size often has a certain degree of impact. On the other hand, when calculating $R(p|c)$ for each patch, we used a fixed-size average pooling kernel to obtain the mean value of the patch. Therefore, we designed an ablation experiment to explore whether different input image sizes and patch sizes significantly affect the performance of the synthetic dataset. In practical applications, we proportionally scale the original image such that its size of shortest edge matches the selected size. For different patch sizes, under the setting of IPC=10, we scale all patches to $112\times 112$ in order to concatenate 4 patches into a $224\times 224$ image. Under the setting of IPC=50, we scale all patches to $224\times 224$.
The experimental results, as shown in Tab. 9, indicate that when the original image size closely matches the standard output size of SD v1.5, better performance are achieved. Additionally, when the size of the cropped patches is closer to the scaled original image size, the image tends to exhibit more complete class-relevant features, leading to better performance. Consider the balance between performance and efficiency, We choose 256x256 as the size of the real images, and $224\times 224$ as the size of the patches. Comparison of visualization with different crop sizes when using 256x256 as the size of the real images are shown in Fig. 9, An excessive size difference between patches and the original image can cause the generated dataset to overly focus on low-frequency features (e.g., noses of dogs), leading to performance degradation.
C Visualizations
To facilitate a more intuitive understanding of our methodology, we present the complete synthetic datasets for ImageWoof and ImageNette at IPC=10 as illustrated in Figs. 10 and 11, as well as the partial synthetic dataset at IPC=50, depicted in Figs. 12 and 13. Furthermore, we have visualized all images belonging to the “Golden Retriever” class at IPC=100, as shown in Fig. 14.